OpenStudy (anonymous):

What is a possible first step in eliminating a variable in the following system of equations? x + 7y = –15 2x + 7y = –9 a. Multiply the first equation by –1. b. Multiply the second equation by –1 c. Multiply the first equation by –2. d. All of the above

OpenStudy (anonymous):

Is it multiply the first equation by -2?

OpenStudy (phi):

The idea is that if you multiply an equation by a number, you can make one of the variables cancel after adding the two equations. multiply by -1 means write -1* in front of every term so a) multiply 1st equation by -1 means -1*x + -1*7y = -1*–15 which simplifies to -x -7y = 15 now add the 2nd equation -x -7y = 15 2x + 7y = –9 2x-x -7y+7y = 15-9 notice the -7y + 7y cancels. so that is a good step.

OpenStudy (anonymous):

oooh, i thought with -2 it would have canceled out as well.

OpenStudy (phi):

However, multiplying the 2nd equation by -1 is also a good step. and if you multiply the 1st equation by -2 you will get a -2x in the first equation, which will cancel +2x in the 2nd so they all look good...

OpenStudy (anonymous):

oh so all of the above.